# Low-Resource Vision Challenges for Foundation Models

### 10 Jan 2024 v2

@@ Abstract
Low-resource settings are well-established in natural language processing, where many languages lack sufficient data for machine learning at scale. However, low-resource problems are under-explored in computer vision. In this paper, we strive to address this gap and explore the challenges of low-resource image tasks with vision foundation models. Thus, we first collect a benchmark of genuinely low-resource image data, covering historic maps, circuit diagrams, and mechanical drawings. These low-resource settings all share the three challenges of data scarcity, fine-grained differences, and the distribution shift from natural images to the specialized domain of interest. While existing foundation models have shown impressive generalizability, we find they cannot transfer well to our low-resource tasks. To begin to tackle the challenges of low-resource vision, we introduce one simple baseline per challenge. Specifically, we propose to i) enlarge the data space by generative models, ii) adopt the best sub-kernels to encode local regions for fine-grained difference discovery and iii) learn attention for specialized domains. Experiments on the three low-resource data sources in our benchmark demonstrate our proposals already provide a better baseline than common transfer learning, data augmentation, and fine-grained methods. This highlights the unique characteristics and challenges of low-resource vision for foundation models that warrant further investigation. Project website: https://xiaobai1217.github.io/Low-Resource-Vision/.
@@ 1 Introduction
As our main contribution, we study the challenges of low-resource vision and collect a low-resource image benchmark. Specifically, our benchmark covers circuit diagrams, historic maps, and mechanical drawings. We find the challenges of low-resource vision tasks are the lack of training data, fine-grained differences, and the domain shift from natural images to specialized domains. From our analysis, we find foundation models struggle to recognize and retrieve low-resource images as do existing transfer learning methods. Thus, we additionally introduce three simple baselines to mitigate each difficulty. Specifically, we finetune foundation models using diverse data produced from generative models to cope with the data scarcity, we discover fine-grained details by focusing on local patterns via selected sub-kernels and we learn attention for specialized domains to combat the distribution shift. Experiments demonstrate the challenges of our low-resource benchmark for existing transfer learning, data augmentation, and fine-grained recognition methods as well as the advantages of our baselines, which can be added to different foundation models. We also discuss the remaining challenges for low-resource vision and the path forward for future works.
@@ 2.1 Tasks
Task I: Circuit Diagram Classification. The goal is to classify the images of circuit diagrams by their function, e.g., audio amplifier, power supply, and signal generator. We manually collect the circuit images and labels from books [18] and websites [1, 2]. In total, we have 32 function classes which we ensure are equally represented in the training set. The challenge of this task comes from small changes in the circuit component giving the circuit a vastly different function. Since there are different layouts for the same function, it is also easy for models to overfit to specific layouts. To measure performance we report Top-1 and Top-5 accuracy.
@@
Table 1: LITE Benchmark Statistics. We show the task formulation and number of images (or image pairs) per split for each of our three tasks. The benchmark will be made publicly available.
@@ 2.2 Low-Resource Vision Challenges
Challenge III: Specialized Domain. Not only is the available data severely limited, but it has a significantly different appearance and comes from an entirely different domain to the natural images commonly used in vision tasks. This means it is both difficult to bootstrap the training data for low-resource tasks with existing datasets and models that are successful on natural images cannot be easily applied to the specialized domains of low-resource images.
@@ 3.1 Baseline I: Generated Data for Data Scarcity
Method. Our proposed baseline is shown in Figure 3. To obtain new images we apply different amounts of Gaussian noise onto a training image $\mathcal{I}$, before applying Stable Diffusion [56] to obtain various augmented images $[\mathcal{I}_{1},\mathcal{I}_{2},\cdots,\mathcal{I}_{m}]$. We denote the amount of noise used to generate $\mathcal{I}_{j}$ as $n_{j}$. With low noise, $n_{j}{<}\gamma$, the generated images are close to the original ones meaning the original label is preserved. With high noise, $n_{j}{>}\tau$, the augmentations become more extreme, often breaking the original label. Since the ground truth for the label-breaking augmentation is unknown we instead propose to utilize such data with self-supervised learning. For this, we adopt the contrastive learning objective [28, 16]. To construct the positive pairs, we generate a second augmented image $\mathcal{I}_{j}^{\prime}$ for each label-breaking augmentation $\mathcal{I}_{j}$ with noise threshold $n_{j}{<}\gamma$. We use these images as a positive pair in a contrastive loss so that the feature of $\mathcal{I}_{j}^{\prime}$ is close to that of the label-breaking image $\mathcal{I}_{j}$, but far away from other label-breaking augmentations. We pass the label-breaking augmentation pairs through the foundation model to obtain their features $x_{j}{=}\mathcal{F}(\mathcal{I}_{j})$ and $x_{j}^{\prime}{=}\mathcal{F}(\mathcal{I}_{j}^{\prime})$, and our objective becomes:
@@ 3.2 Baseline II: Tokenization for Fine-Grained
To obtain output features with the same dimension as in the original projection, we apply max pooling $\mathbf{b}^{\prime}{=}\text{MaxPool}(\mathbf{b})$ and flatten $\mathbf{b^{\prime}}$. This allows us to use the existing positional encodings and class tokens and ensures our input is suitable for the frozen layers of the foundation model. Since only the weighting $\mathbf{w}\in\mathbb{R}^{T}$ is learned, our tokenization can be easily trained on low-resource data and effectively encourages the foundation model to focus on fine-grained details.
@@ 3.3 Baseline III: Attention for Specialized Domains
where $\alpha$ is a learnable parameter to balance the two attentions. As the only learnable parameters are $\bm{\mathcal{M}}$, $\alpha$, and $\mathbf{r}$, training our attention for specialized domains allows adaptation without overfitting.
@@ 4 Related Work
High-Resource Vision. The large majority of computer vision research focuses on high-resource settings, where data is plentiful. Various benchmarks of high-resource images have been proposed, e.g., [57, 49, 17, 24, 53, 52, 45], unlocking the ability to train larger and larger models. Their images are crawled from the internet [57, 49, 52, 45], or captured by the authors [24, 17, 53]. The labels can be either coarse-grained [57, 49, 24, 17, 77, 10] or fine-grained [52, 45, 53, 75, 40]. High-resource vision tends to focus mostly on natural images which are the most plentiful online. However, some benchmarks also collect images from other domains, e.g., X-ray [65], underwater [32], medical [36] and satellite [31]. These are less high-resource than natural images. However, they still contain at least several thousands of samples per benchmark. Different from previous high-resource image datasets, we focus on the low-resource settings, where images are severely limited with only a few hundred samples available for training.
@@
Vision Foundation Models. Vision foundation models are pre-trained by high-resource web-crawled images with weak supervision (like captions) or human annotations, and present impressive generalizability on various downstream tasks. While CLIP [55], BLIP [47] and ALIGN [38] learn from image-text pairs only, ImageBind [25] uses image-paired data of multiple modalities. The recent SAM [71] and UniDetector [42] instead propose foundation models for object detection and segmentation. However, the impressive generalization ability has been focused on natural images, likely similar to many in the large-scale training set [74]. Simultaneously, there are many low-resource problems from specialized image domains lacking a large amount of online data, such as technical images. In this paper, we create a benchmark of low-resource vision problems and demonstrate that foundation models cannot generalize to such data.
@@
It is also possible to adapt the strong image representation learned by foundational models to new tasks. This can be simply done by model fintuning [76] or by training an additional projection layer [23] on the pre-trained model features. Several works [34, 15, 26, 50, 58] instead add new parameters into the layers of a pre-trained model, only training these parameters. Although these works achieve impressive performance, they are not suited for low-resource vision where the training data is severely limited and from fine-grained, specialized domains that are highly dissimilar to the pre-training data. We propose to study such tasks and their challenges and propose baselines for better adaptation to low-resource tasks.
@@
Low-Shot Vision. A huge number of works have studied scenarios with limited training data. One typical setting is few-shot learning [64, 46, 12], where the goal is to generalize to previously unseen classes with only a few samples available for model learning. Some works study the related task of in-context learning [81, 9, 70, 61, 60], which allows inference on unseen tasks by conditioning on related examples without updating the model parameters. Other works reduce this one step further and study zero-shot scenarios [8, 72, 73, 80, 14, 43], where no data of the relevant classes are seen in training, although prior knowledge or data from other classes could be used. All these works make a significant step towards reducing the amount of data needed for model learning. However, none of these tasks study the combination of scarce data, fine-grained differences and highly specialized domains present in low-resource vision.
@@ 5.1 Difficulties for Vision Foundation Models
To understand how well current vision foundation models address low-resource vision tasks, we first examine their zero-shot performance on our LITE benchmark. We consider four vision foundation models that are publicly available: CLIP [55], BLIP [47], SAM [42] and ImageBind [25].
Setup. To obtain zero-shot results for circuit diagram classification, we follow [55] and customize the label text to make it better suited to the models. Specifically, we use the prompt template “A circuit diagram of {label}.”, where the label is a category label, e.g., power supply or motor driver. We cannot obtain zero-shot results for SAM [42] in this way, so we omit it for circuit diagram classification. For all tasks, we simply calculate the similarities among the feature embeddings between the input image and the ground-truth image or text to find the closest neighbors.
Results. From the results in Table 2 we observe none of these foundation models perform well on low-resource settings. Although ImageBind is better suited due to its larger pre-training set size and image-focused embedding, there is still much room for improvement in low-resource vision. Despite current foundation models’ impressive generalizability on other benchmarks, they cannot yet solve the combined challenges of data scarcity, fine-grained details, and highly specialized domains. Unlike other zero-shot and few-shot tasks where foundation models have shown good generalization, low-resource data is truly scarce online. This means it is unlikely to be present in the training data of foundation models. The specialized domain means it is also highly dissimilar to natural images which form a large part of foundation model training data [74]. Due to the model’s unfamiliarity with low-resource data, it struggles to attend to fine-grained, task-relevant details. Therefore, vision foundation models need adaptation for low-resource tasks.
@@
Table 4: Combination with Different Foundation Models. Our approach can easily be plugged into different foundational models, improving their ability to adapt to low-resource tasks. However, the tasks are still far from solved demonstrating the need for further study of low-resource vision.
Circuit Classification
Top-1 (%) $\uparrow$ Top-5 (%) $\uparrow$
Zero-Shot Transfer 19.3 45.1
Fine-Grained
Adaptive-FGSBIR [13] 16.7 43.2
PLEor [68] 17.1 44.1
PDiscoNet [63] 16.2 43.5
Our Baselines
Tokenization for Fine-Grained 20.9 45.5
Combination of Baselines 24.1 49.3
5.2 Challenge Results
Setup. As ImageBind obtains the best zero-shot transfer performance on our low-resource benchmark, we adapt this model by the existing transfer learning method AdaptFormer [15] in the following experiments. For all three challenges defined in Section 2 we add our proposed baselines or existing methods to this default setting, keeping the foundation model parameters frozen. Our baselines work independently of each other, focusing on different areas of the foundation model: input, tokenization, and attention. Thus they can be easily combined. We test this combination as well as the individual baselines.
Challenge I: Data Scarcity. In Table 3, we demonstrate the challenge of low-resource vision for existing solutions to data scarcity. Specifically, we test popular data augmentation methods on circuit diagram classification. Traditional data augmentation methods like random cropping and flipping as well as CutMix [78] struggle with our low-resource benchmark. When using such a small set of images with very fine-grained differences these methods deliver limited improvements to the data diversity. Mixup [79] and SyntheticData [29] obtain better performance as they are able to create more diverse training data by mixing samples and utilizing generative models. Although DA-Fusion [62] and SyntheticData [29] use generative models to obtain more data points, they only consider generated images that are similar to the original ones, i.e. label-preserving. In contrast, our baseline considers both label-preserving and label-breaking generated images and can therefore benefit from many more data points crucial for low-resource vision tasks. This is demonstrated further in the appendix with the results for the other two tasks. Combining our baseline solutions to all three challenges results in the best performance, highlighting the multifaceted nature of low-resource vision.
@@ 5.3 Our Baselines on Different Foundation Models
In Table 4, we demonstrate that our low-resource baselines can be plugged into different foundational models by adding them to CLIP [55], BLIP [47], SAM [42] and ImageBind [25]. We observe that all four foundation models can be improved by a large margin with adaptation to low-resource tasks. For example, by adding AdaptFormer, we observe +33.7% R@1 for CLIP and +41.1% R@1 for ImageBind on mechanical drawing retrieval. The adaptation allows the foundation model features to be better suited to the specific domain of a low-resource task and can thus distinguish images with distinctive patterns. Adding our simple baselines results in further improvements. For example, there is an additional +4.7% R@1 improvement for CLIP and +5.7% for ImageBind on mechanical drawing retrieval. Nevertheless, the performance of our baselines with the best-performing foundation model is still low, e.g., 24.1% Top-1 accuracy on circuit diagram classification and 37.1 R@1 on historic map retrieval. Thus, the proposed low-resource tasks are still far from solved and warrant further study.

### 11 Apr 2024 v3

@@ Abstract
Low-resource settings are well-established in natural language processing, where many languages lack sufficient data for deep learning at scale. However, low-resource problems are under-explored in computer vision. In this paper, we address this gap and explore the challenges of low-resource image tasks with vision foundation models. We first collect a benchmark of genuinely low-resource image data, covering historic maps, circuit diagrams, and mechanical drawings. These low-resource settings all share three challenges: data scarcity, fine-grained differences, and the distribution shift from natural images to the specialized domain of interest. While existing foundation models have shown impressive generalizability, we find they cannot transfer well to our low-resource tasks. To begin to tackle the challenges of low-resource vision, we introduce one simple baseline per challenge. Specifically, we i) enlarge the data space by generative models, ii) adopt the best sub-kernels to encode local regions for fine-grained difference discovery and iii) learn attention for specialized domains. Experiments on our three low-resource tasks demonstrate our proposals already provide a better baseline than transfer learning, data augmentation, and fine-grained methods. This highlights the unique characteristics and challenges of low-resource vision for foundation models that warrant further investigation. Project page: https://xiaobai1217.github.io/Low-Resource-Vision/.
@@ 1 Introduction
As our main contribution, we study the challenges of low-resource vision and collect a low-resource image benchmark. Specifically, our benchmark covers circuit diagrams, historic maps, and mechanical drawings. We find the challenges of low-resource vision are a lack of training data, fine-grained differences, and domain shift from natural images to specialized domains. From our analysis, we discover foundation models struggle to recognize and retrieve low-resource images as do existing transfer learning methods. Thus, we introduce three simple baselines to mitigate each difficulty. Specifically, we finetune foundation models using diverse data produced by generative models to cope with data scarcity, we discover fine-grained details by focusing on local patterns via selected sub-kernels, and we learn attention for specialized domains to combat the distribution shift. Experiments demonstrate the challenges of our low-resource benchmark for existing transfer learning, data augmentation, and fine-grained methods as well as the advantages of our baselines, which can be added to different foundation models. We also discuss the remaining challenges for low-resource vision and paths forward for future works.
@@ 2.1 Tasks
Task I: Circuit Diagram Classification. The goal is to classify the images of circuit diagrams by their function, e.g., audio amplifier and power supply. We collect circuit images and labels from books [19] and websites [1, 2, 3]. In total, we have 32 function classes which are equally represented in training. The challenge comes from small changes in circuit components dramatically changing the function. Since there are different layouts for the same function, it is also easy for models to overfit to specific layouts. We measure performance with Top-1 and Top-5 accuracy.
@@
Table 1: LITE Benchmark Statistics. We show the task formulation and number of images (or image pairs) per split for each of our three tasks. The benchmark is available on our project website.
@@ 2.2 Low-Resource Vision Challenges
Challenge III: Specialized Domain. Not only is the available data severely limited, but it has a significantly different appearance to the natural images commonly used in vision tasks. This means it is difficult to bootstrap the training data for low-resource tasks with existing datasets. Moreover, models that are successful on natural images cannot be easily applied to the specialized domains of low-resource images.
@@ 3.1 Baseline I: Generated Data for Data Scarcity
Method. Our proposed baseline is shown in Figure 3. In Stable Diffusion [59], the forward process gradually adds Gaussian noise to an image with a variance schedule $\beta_{1},...,\beta_{T}$ ($T{=}50$). To obtain new images, we sample noisy images at different timestep $t$ and start the reverse process. For label-preserving augmentations we want a small $t$ so the generated image is close to the original and thus adopt $\gamma{=}0.3$ and $t{=}\gamma\cdot T$. For label-breaking augmentations we use $\tau{=}0.6$ and $t{=}\tau\cdot T$. Then, we can obtain various augmented images $[\mathcal{I}_{1},\mathcal{I}_{2},\cdots,\mathcal{I}_{m}]$. Since the ground truth for the label-breaking augmentation is unknown we instead utilize such data with a contrastive learning objective [30, 17]. To construct the positive pairs, we generate a second augmented image $\mathcal{I}_{j}^{\prime}$ for each label-breaking augmentation $\mathcal{I}_{j}$ with sampling timestep $t{=}\gamma\cdot T$. The contrastive loss encourages the feature of $\mathcal{I}_{j}^{\prime}$ to be close to that of the label-breaking image $\mathcal{I}_{j}$, but far away from other label-breaking augmentations. We pass the label-breaking augmentation pairs through the foundation model to obtain their features $x_{j}{=}\mathcal{F}(\mathcal{I}_{j})$ and $x_{j}^{\prime}{=}\mathcal{F}(\mathcal{I}_{j}^{\prime})$, and our objective becomes:
@@ 3.2 Baseline II: Tokenization for Fine-Grained
To obtain output features with the same dimension as the original projection, we apply max pooling $\mathbf{b}^{\prime}{=}\text{MaxPool}(\mathbf{b})$ and flatten $\mathbf{b^{\prime}}$. We can then use the existing positional encodings and class tokens and ensure our input is suitable for the frozen foundation model. As only the weighting $\mathbf{w}\in\mathbb{R}^{T}$ is learned, our tokenization is suitable for low-resource data and effectively encourages focus on fine-grained details.
@@ 3.3 Baseline III: Attention for Specialized Domains
where $\alpha$ is learned to balance the two attentions. As only $\bm{\mathcal{M}}$, $\alpha$, and $\mathbf{r}$ are learned, training our attention for specialized domains allows adaptation without overfitting.
@@ 4 Related Work
High-Resource Vision. The large majority of computer vision research focuses on high-resource settings, where data is plentiful. Various benchmarks of high-resource images have been proposed, [60, 51, 18, 26, 55, 54, 47], unlocking the ability to train larger and larger models. Their images are crawled from the internet [60, 51, 54, 47], or captured by the authors [26, 18, 55]. The labels can be either coarse-grained [60, 51, 26, 18, 80, 11] or fine-grained [54, 47, 55, 78, 42]. High-resource vision tends to focus on natural images which are plentiful online. However, some benchmarks also collect images from other domains, e.g., X-ray [68], underwater [34], medical [38] and satellite [33]. These are less high-resource than natural images. However, they still contain thousands of samples. Different from previous high-resource image datasets, we focus on low-resource settings, where images are severely limited with only a few hundred samples available for training.
@@
Vision Foundation Models. Vision foundation models are pre-trained by high-resource web-crawled images with weak supervision or human annotations, and present impressive generalizability on various downstream tasks. While CLIP [58], BLIP [49], and ALIGN [40] learn from image-text pairs only, ImageBind [27] uses image-paired data of multiple modalities. Recent works SAM [74], DINOv2 [56], UniDetector [44] and AIM [22] instead propose foundation models for visual-only tasks such as object detection, segmentation and depth estimation. However, the impressive generalization ability has been focused on natural images, likely similar to many in the large-scale training set [77]. Simultaneously, there are many low-resource problems from specialized domains lacking a large amount of online data, such as technical images. In this paper, we create a benchmark of low-resource vision problems and demonstrate that foundation models cannot generalize to such data.
@@
It is also possible to adapt the strong image representations of foundational models to new tasks. This can be done by fintuning [79] or by training additional projection layers [25]. Several works [36, 16, 28, 52, 61] instead add new trainable parameters into the layers of a frozen pre-trained model. Although these works achieve impressive performance, they are not suited for low-resource vision where training data is severely limited and from fine-grained, specialized domains that are highly dissimilar to the pre-training data. We study such tasks and their challenges and propose baselines for better adaptation to low-resource tasks.
@@
Low-Shot Vision. A huge number of works have studied scenarios with limited training data. One typical setting is few-shot learning [67, 48, 13], which aims to generalize to previously unseen classes with only a few training samples. Some works study in-context learning [84, 10, 73, 64, 63], which allows inference on unseen tasks by conditioning on related examples without updating the model parameters. Other works reduce this one step further and study zero-shot scenarios [9, 75, 76, 83, 15, 45], where no data of the relevant classes are seen in training, although prior knowledge or data from other classes could be used. All these works make a significant step towards reducing the amount of data needed for model learning. However, none of these tasks study the combination of scarce data, fine-grained differences and highly specialized domains present in low-resource vision.
@@ 5.1 Difficulties for Vision Foundation Models
To understand how well current vision foundation models address low-resource vision tasks, we first examine their zero-shot performance on our LITE benchmark. We consider six vision foundation models: CLIP [58], BLIP [49], SAM [44], AIM [22], DINOv2 [56] and ImageBind [27].
Setup. To obtain zero-shot results for circuit diagram classification, we follow [58] and customize the label text to make it better suited to the models. Specifically, we use the prompt template “A circuit diagram of {label}.”, where the label is a category label, e.g., power supply or motor driver. We cannot obtain zero-shot results for SAM, AIM, and DINOv2 in this way, so we omit these models for circuit diagram classification. For all tasks, we calculate the similarities among the feature embeddings between the input image and the ground-truth image or text to find the closest neighbors.
Results. From Table 2 we observe none of these foundation models perform well on low-resource tasks. Although ImageBind is better suited due to its larger pre-training set and image-focused embedding, there is still much room for improvement. Despite current foundation models’ impressive generalizability on other benchmarks, they cannot yet solve the combined challenges of data scarcity, fine-grained details, and highly specialized domains. Unlike other zero-shot and few-shot tasks where foundation models have shown good generalization, low-resource data is truly scarce online, meaning it is unlikely to be in the training data of foundation models. The specialized domain means it is also highly dissimilar to natural images which form a large part of foundation model training data [77]. Due to the models’ unfamiliarity with low-resource data, they struggle to attend to fine-grained, task-relevant details. Therefore, vision foundation models need adaptation for low-resource tasks.
@@
5.2 Challenge Results
Setup. As ImageBind obtains the best zero-shot performance on our low-resource benchmark, we use it in this section. For all three challenges defined in Section 2 we add our proposed baselines or existing methods alongside AdaptFormer [16], keeping the foundation model frozen. Our baselines are independent of each other, focusing on different areas of the foundation model: input, tokenization, and attention. Thus they can be easily combined. We test this combination as well as the individual baselines.
Circuit Classification
Top-1 (%) $\uparrow$ Top-5 (%) $\uparrow$
Zero-Shot Transfer 19.3 45.1
Fine-Grained
Adaptive-FGSBIR [14] 16.7 43.2
PLEor [71] 17.1 44.1
PDiscoNet [66] 16.2 43.5
Our Baselines
Tokenization for Fine-Grained 20.9 45.5
Combination of Baselines 24.1 49.3
@@ 5.2 Challenge Results
AIM [22]
Zero-Shot Transfer - - 12.0 33.0 37.9 14.9 31.2 72.2
AdaptFormer 16.3 41.8 16.4 37.8 32.5 55.2 78.3 12.7
Our Baselines 20.1 45.7 20.3 41.2 28.8 59.4 82.9 10.0
DINOv2 [56]
Zero-Shot Transfer - - 1.5 7.1 83.4 15.9 32.2 83.0
AdaptFormer 15.8 40.3 13.6 16.9 58.5 56.0 79.1 16.5
Our Baselines 20.3 45.8 18.2 21.9 54.1 60.7 83.1 12.4
@@
Table 5: Combination with Different Foundation Models. Our approach can easily be applied to different foundational models, improving their adaptation to low-resource tasks. However, the tasks are far from solved highlighting the need for further study of low-resource vision.
Challenge I: Data Scarcity. Table 3 demonstrates the challenge of low-resource vision for existing solutions to data scarcity. Specifically, we test popular data augmentation methods on circuit classification. Traditional methods like random crop and flip as well as CutMix [81] struggle with our LITE benchmark. When using such a small set of images with very fine-grained differences these methods deliver limited additional data diversity. Mixup [82] and SyntheticData [31] obtain better performance as they can create more diverse training data by mixing samples and utilizing generative models. Although DA-Fusion [65] and SyntheticData [31] use generative models to obtain more data, they only consider generated images that are similar to the original ones, i.e. label-preserving. In contrast, our baseline considers both label-preserving and label-breaking generated images and therefore benefits from many more data points crucial for low-resource vision tasks. This is demonstrated further in the appendix with results for the other two tasks. Combining our baseline solutions to all three challenges results in the best performance, highlighting the multifaceted nature of low-resource vision.
@@ 5.3 Our Baselines on Different Foundation Models
In Table 5, we demonstrate that our low-resource baselines can be plugged into different foundational models by adding them to CLIP [58], BLIP [49], SAM [44], AIM [22], DINOv2 [56] and ImageBind [27]. All six foundation models can be improved by a large margin with adaptation to low-resource tasks. For example, by adding AdaptFormer, we observe +33.7% R@1 for CLIP, +40.3% R@1 for AIM, +40.1% R@1 for DINOv2 and +41.1% R@1 for ImageBind on mechanical drawing retrieval. The adaptation allows the foundation model features to be better suited to the specific domain of a low-resource task and can thus distinguish images with distinctive patterns. Adding our simple baselines results in further improvements. For example, there is an additional +4.7% R@1 improvement for CLIP and DINOv2 and +5.7% for ImageBind on mechanical drawing retrieval. Nevertheless, the performance of our baselines with the best-performing foundation model is still low, e.g., 24.1% Top-1 accuracy on circuit diagram classification and 37.1 R@1 on historic map retrieval. Thus, the proposed low-resource tasks are still far from solved and warrant further study.